# GS-Occ3D: Scaling Vision-only Occupancy Reconstruction for Autonomous Driving with Gaussian Splatting

### v1

@@ 4.5 Ablation Studies
Input Camera Views. As shown in Tab.˜3, compared to using only 3 cameras, our method achieves better geometry reconstruction when utilizing the full panoramic 5-camera input. In contrast, some existing methods show reduced reconstruction and rendering quality with 5-camera input compared to 3-camera input due to the added geometry and rendering ambiguities introduced by forward-facing multi-view inputs in street scenes. However, our octree-based representation with LoD maintains structural fidelity and consistency across both global and local scales, enabling effective exploitation of multi-view observations. This makes our method well-suited for reconstructing diverse crowdsourced data and producing high-quality label curation.
@@
Effectiveness of Our Ground Gaussians. Fig.˜8 demonstrates the superiority of our ground gaussians in capturing detailed features of weakly-textured areas. Without separate handling of the ground, holes and abnormal protrusions can occur, distorting the geometry.
@@ 5 Conclusion
GS-Occ3D is a scalable vision-only occupancy reconstruction framework that uses crowdsourced data from consumer vehicles for auto-labeling. It enables cost-efficient and scalable occupancy label curation. First, we adopt an Octree-based Gaussian Surfel formulation to reconstruct geometry for ground, background, and dynamic objects. Our vision-only reconstruction achieves SOTA geometric results. We then reconstruct the entire Waymo dataset, which covers diverse scenes. This enables us to validate the effectiveness of our labels for downstream occupancy models on Occ3D-Waymo and demonstrate superior zero-shot generalization on Occ3D-nuScenes. Our results highlight the potential of large-scale vision-based occupancy reconstruction as a new paradigm for autonomous driving perception. We hope this work inspires further research into using reconstruction to empower vision-only label curation for downstream autonomous driving tasks.
Limitations. Our method has several limitations. (1) Cameras inherently provide only front and side views, lacking rear view coverage, which leads to unavoidable information loss. (2) Under nighttime conditions or exposure anomalies, the effective observation range of vision-based methods is significantly reduced. (3) In ego-static scenarios, vision-only methods often fail to reconstruct geometry. Methods such as [80, 35, 18], which rely on pretraining with Waymo LiDAR point clouds, are incompatible with the vision-only setting. More recent methods such as [76, 13] may overcome it. (4) We specifically focused on geometry reconstruction, consequently providing results solely on geometry generalization. Future work will aim to enhance both semantic and geometry reconstruction, particularly under challenging weather and lighting conditions, to achieve greater robustness in real-world applications.

### v3

@@ 4.5 Ablation Studies
Input Camera Views. As shown in Tab.˜3, our method improves reconstruction with a 5-camera input over 3. In contrast, the performance degradation in other methods is primarily caused by the geometric and rendering ambiguities from forward-facing multi-view inputs. However, our octree-based representation with LoD maintains structural fidelity and consistency across both global and local scales, enabling effective exploitation of multi-view observations. This makes our method well-suited for reconstructing diverse crowdsourced data and generating high-quality label.
@@
Effectiveness of Our Ground Gaussians. Fig.˜8 shows the effectiveness in capturing geometry of weakly-textured areas. Without separate handling of the ground, holes and abnormal protrusions can occur, distorting the geometry.
@@ 5 Conclusion
GS-Occ3D is a scalable vision-only occupancy reconstruction framework that uses crowdsourced data from consumer vehicles for auto-labeling. It enables cost-efficient and scalable occupancy label curation. First, we adopt an Octree-based Gaussian Surfel formulation to reconstruct geometry for ground, background, and dynamic objects. Our vision-only reconstruction achieves SOTA geometric results. We then reconstruct the entire Waymo dataset, which covers diverse scenes. This enables us to validate the effectiveness of our labels for downstream occupancy models on Occ3D-Waymo and demonstrate superior zero-shot generalization on Occ3D-nuScenes. Our results highlight the potential of large-scale vision-based occupancy reconstruction as a new paradigm for scalable auto-labeling. We hope this work inspires further research into using advanced reconstruction to empower vision-only label curation for downstream tasks.
Limitations. (1) Cameras inherently provide only front and side views, lacking rear view coverage, which leads to unavoidable information loss. (2) Performance degrades at night or with exposure issues, which reduce the effective visual range. (3) Vision methods often fail in ego-static scene. While methods pretrained on Waymo LiDAR [80, 35, 18] are incompatible with our vision-only setting, recent works [76, 13] may offer solutions. (4) We focus on geometry reconstruction, providing results solely on geometry generalization. Future work will aim to enhance both semantic and geometry reconstruction, particularly under challenging weather and lighting conditions, to achieve greater robustness in real-world applications.
Acknowledgments
This work is supported by the National Key Research and Development Program of China (2022YFB4701900), and the Tsinghua University - Mercedes Benz Joint Institute for Sustainable Mobility.